What is the formula or method of coming up with a line that is perpendicular to and equation and passes through a specific point...
ex: equation of a line that is perpendicular to 3x-4y=10 and passes through (-1,2)

Answer by scott8148(6628)

perpendicular lines have slopes which are negative-reciprocals of each other
eg. 3/2 and -2/3

when the equation is in standard form (as in this problem) ; swapping the x and y coefficients
and changing one of the signs
will give the perpendicular line

to have the line go through a particular point , plug the coordinates into the equation
to get a new constant term

4x + 3y = 10 is perpendicular to 3x - 4y = 10

4(-1) + 3(2) = 2 ___ 4x + 3y = 2 goes through (-1,2)